# Training-Free Adaptive Quantization for Variable Rate Image Coding for Machines

### v1

@@ Abstract
Image Coding for Machines (ICM) has become increasingly important with the rapid integration of computer vision into real-world applications. However, most ICM frameworks utilize learned image compression (LIC) models that operate at a fixed rate and require separate training for each target bitrate, which may limit their practical applications. Existing variable rate LIC approaches mitigate this limitation but typically depend on training, increasing computational cost and deployment complexity. Moreover, variable rate control has not been thoroughly explored for ICM. To address these challenges, we propose a training-free, adaptive quantization step size control scheme that enables flexible bitrate adjustment. By leveraging both channel-wise entropy dependencies and spatial scale parameters predicted by the hyperprior network, the proposed method preserves semantically important regions while coarsely quantizing less critical areas. The bitrate can be continuously controlled through a single parameter. Experimental results demonstrate the effectiveness of our proposed method, achieving up to 11.07% BD-rate savings over the non-adaptive variable rate method.
@@ I Introduction
In this paper, we propose a training-free variable rate framework for ICM. Our method enables flexible bitrate adjustment by dynamically modulating the quantization step size according to both channel-wise and spatial characteristics of the latent representation. Consequently, it preserves semantically important regions while coarsely quantizing redundant areas. The proposed design further allows continuous bitrate control with a single parameter without training or additional side information. Owing to its simplicity and generality, our method can be readily integrated into existing ICM architectures for practical deployment.
@@ II Related Work
Figure 1: Visualization of average scale parameters of each (a) channel and (b) slice. Deeper colors represent larger values.
@@ III-A Preliminary Findings
To investigate the channel-wise characteristics of the scale parameters predicted by the entropy model, we conduct a preliminary experiment using an image from the Kodak dataset [35]. The image is compressed with SA-ICM, and the average scale values across channels and slices are visualized in Fig. 1. The results reveal that channels belonging to earlier slices tend to have larger scale values, which indicates the greater bit allocation, whereas those in later slices contain more redundant information. This trend is consistent with the observations reported by D. He et al. [16], although their study utilized human-oriented LIC.
We also examine the spatial characteristics of the scale parameters, as illustrated in Fig. 2. Larger scale values are observed along object boundaries, while smaller values appear in other regions such as object interiors and backgrounds. These results indicate that the scale parameter reflects region-wise importance relevant to recognition tasks within the image. Consequently, adapting the quantization step size according to the scale can assign finer quantization to important regions and coarser quantization to less informative areas. This observation forms the key intuition behind our proposed method.
@@ III-B Overview of Proposed Quantization Controller
In this paper, we propose a training-free variable rate method for machine-oriented image compression. The compression process is illustrated in Fig. 3. In contrast to non-adaptive step size control, which uniformly scales the quantization step size across all features, our approach adaptively adjusts it. Specifically, the quantization controller first decides the maximum and minimum quantization step sizes for each channel using channel-wise dependencies captured by the autoregressive entropy model. Then, the actual step size for each latent pixel is determined within this range using spatial scale variations predicted by the hyperprior network. Note that the bitrate can be continuously controlled by a single parameter $d$, which facilitates deployment in real-world scenarios. The $d$ value is defined for $d>0$; as $d$ increases, the transmitted bitrate decreases. When $d=1$, the bitrate matches that of the base model. While the quantization step size of the latent representation $\boldsymbol{y}$ is adaptively varied for rate control, that of the hyper latent $\boldsymbol{z}$ is fixed at 1 because coarse quantization of $\boldsymbol{z}$ would impair its role in predicting the variance of $\boldsymbol{y}$. Moreover, since the bitrate contribution of $\boldsymbol{z}$ is relatively small [27], its variation has negligible impact on overall rate control.
@@ III-C Channel-Aware Quantization Step Size Control
The preliminary experiments revealed that the scale parameter varies across channels. However, in Ch-ARM, the exact scales for all slices are not available simultaneously, making direct scheduling based on the true scales infeasible. Motivated by the empirical observation that earlier slices carry more critical information while later slices are increasingly redundant, we assign larger quantization step sizes to channels in later slices. Specifically, the user-specified parameter $d$ determines the quantization step size bounds. When $0<d<1$, the minimum quantization step size for the $n$-th slice is linearly decreased, while when $d>1$, the maximum size is linearly increased. The minimum and maximum quantization step size of the $n$-th slice, $\Delta_{\min}^{(n)}$ and $\Delta_{\max}^{(n)}$, are defined as:
@@ V Conclusion
In this paper, we propose a training-free quantization controller for variable rate ICM that adaptively adjusts quantization step size based on channel-wise and spatial characteristics. By assigning larger quantization steps to later channels and redundant regions according to the scale parameters predicted by the entropy model, our method effectively reduces bitrate without training. Additionally, the bitrate can be continuously controlled by a single parameter. Experimental results demonstrate that our proposed method outperforms a variable rate baseline with non-adaptive quantization. Future work will further extend this framework to scalable image coding for humans and machines.

### v2

@@ Abstract
Image Coding for Machines (ICM) has become increasingly important with the rapid integration of computer vision technology into real-world applications. However, most neural network-based ICM frameworks operate at a fixed rate, thus requiring individual training for each target bitrate. This limitation may restrict their practical usage. Existing variable rate image compression approaches mitigate this issue but often rely on additional training, which increases computational costs and complicates deployment. Moreover, variable rate control has not been thoroughly explored for ICM. To address these challenges, we propose a training-free quantization strength control scheme that enables flexible bitrate adjustment. By exploiting the scale parameter predicted by the hyperprior network, the proposed method adaptively modulates quantization step sizes across both channel and spatial dimensions. This allows the model to preserve semantically important regions while coarsely quantizing less critical areas. Our architectural design further enables continuous bitrate control through a single parameter. Experimental results demonstrate the effectiveness of our proposed method, achieving up to 11.07% BD-rate savings over the non-adaptive variable rate baseline.
@@ I Introduction
In this paper, we propose a training-free variable rate framework for ICM. Our method enables flexible bitrate adjustment by dynamically controlling the quantization strength according to both channel-wise and spatial characteristics of the latent representation. Consequently, it preserves semantically important regions while coarsely quantizing redundant areas. The proposed design further allows continuous bitrate control with a single parameter, without requiring additional training or side information. Owing to its simplicity and generality, our method can be readily integrated into existing ICM architectures for practical deployment.
@@ II Related Work
Figure 1: Visualization of average scale values of each (a) channel and (b) slice. Deeper colors represent larger values.
@@ III-A Preliminary Findings
Since the scale parameter predicted by the hyperprior network determines bit allocation, we hypothesize that in ICM, it correlates with the feature importance required for recognition. To investigate the channel-wise characteristics of the scale values, we conduct a preliminary experiment using an image from the Kodak dataset [35]. The image is compressed with SA-ICM, and the average scale values across channels and slices are visualized in Fig. 1. The results reveal that channels belonging to earlier slices tend to have larger scale values, which indicates the greater bit allocation, whereas those in later slices contain more redundant information. This trend is consistent with the observations reported by D. He et al. [16], although their study utilized human-oriented LIC.
We also examine the spatial characteristics of the scale parameters, as illustrated in Fig. 2. Larger scale values are observed along object boundaries, while smaller values appear in other regions such as object interiors and backgrounds. Given that SA-ICM is designed to preserve edge information effectively, these results indicate that the scale parameter reflects the region-wise importance relevant to recognition tasks. This observation forms the key intuition behind our proposed method.
@@ III-B Overview of Proposed Quantization Controller
In this paper, we propose a training-free variable rate method for ICM. The compression process is illustrated in Fig. 3. Since our preliminary findings reveal that the scale parameter serves as an indicator of feature importance for recognition, our approach adaptively adjusts the quantization strength according to this parameter to achieve variable bitrates without degrading recognition accuracy. Specifically, the quantization controller first decides a channel-wise range of allowable quantization step sizes by utilizing scale characteristics. The actual step size for each latent pixel is then determined within this range based on spatial scale variations. Note that the bitrate can be continuously controlled by a single parameter $d$, which facilitates deployment in real-world scenarios. The $d$ value is defined for $d>0$; as $d$ increases, the transmitted bitrate decreases. When $d=1$, the bitrate matches that of the base model. While the quantization step size of the latent representation $\boldsymbol{y}$ is adaptively varied for rate control, that of the hyper latent $\boldsymbol{z}$ is fixed at 1 because coarse quantization of $\boldsymbol{z}$ would impair its role in predicting the variance of $\boldsymbol{y}$. Moreover, since the bitrate contribution of $\boldsymbol{z}$ is relatively small [27], its variation has negligible impact on overall rate control.
@@ III-C Channel-Aware Quantization Step Size Control
The preliminary experiments reveals that the scale parameter varies across channels. However, in Ch-ARM, the exact scales for all slices are not available simultaneously, making direct scheduling based on the true scales infeasible. Motivated by the empirical observation that earlier slices carry more critical information while later slices are increasingly redundant, we assign larger quantization step sizes to channels in later slices. Specifically, the user-specified parameter $d$ determines the quantization step size bounds. When $0<d<1$, the minimum quantization step size for the $n$-th slice is linearly decreased, while when $d>1$, the maximum size is linearly increased. The minimum and maximum quantization step size of the $n$-th slice, $\Delta_{\min}^{(n)}$ and $\Delta_{\max}^{(n)}$, are defined as:
@@ V Conclusion
In this paper, we propose a training-free quantization controller for variable rate ICM that adaptively adjusts quantization step size based on channel-wise and spatial characteristics. By assigning larger quantization steps to later channels and redundant regions according to the scale parameter predicted by the hyperprior network, our method effectively reduces bitrate without training. Additionally, the bitrate can be continuously controlled by a single parameter. Experimental results demonstrate that our proposed method outperforms a variable rate baseline with non-adaptive quantization. Future work will further extend this framework to scalable image coding for humans and machines.